# Human Mesh Modeling for Anny Body

### v1

@@ Abstract
Parametric body models are central to many human-centric tasks, yet existing models often rely on costly 3D scans and learned shape spaces that are proprietary and demographically narrow. We introduce Anny, a simple, fully differentiable, and scan-free human body model grounded in anthropometric knowledge from the MakeHuman community. Anny defines a continuous, interpretable shape space, where phenotype parameters (e.g. gender, age, height, weight) control blendshapes spanning a wide range of human forms — across ages (from infants to elders), body types, and proportions. Calibrated using WHO population statistics, it provides realistic and demographically grounded human shape variation within a single unified model. Thanks to its openness and semantic control, Anny serves as a versatile foundation for 3D human modeling — supporting millimeter-accurate scan fitting, controlled synthetic data generation, and Human Mesh Recovery (HMR). We further introduce Anny-One, a collection of 800k photorealistic humans generated with Anny, showing that despite its simplicity, HMR models trained with Anny can match the performance of those trained with scan-based body models, while remaining interpretable and broadly representative. The Anny body model and its code are released under the Apache 2.0 license, making Anny an accessible foundation for human-centric 3D modeling.
@@ 1 Introduction
Understanding human shape and motion in 3D is fundamental to computer vision, graphics, and robotics. Parametric body models provide a compact and differentiable representation of the human body, enabling a wide range of applications — from scan fitting and animation to Human Mesh Recovery (HMR) from images and videos [goel2023humans, yin2025smplest, zhang2021pymaf, shin2024wham, kocabas2020vibe]. These models have been central to progress in human-centric perception, powering tasks such as motion analysis, behavior understanding, and human–robot interaction.
Among them, the SMPL family [loper2015smpl, pavlakos2019smplx, osman2020star, osman2022supr] has played a transformative role. SMPL-like models achieve high accuracy on standard benchmarks by learning low-dimensional shape spaces from 3D body scans [robinette2002caesar, Bogo:CVPR:2014, Bulat2017FAW]. However, this data-driven design also imposes limitations: the available scan datasets are expensive, demographically narrow, and privacy-sensitive. As a result, current models struggle to represent the full diversity of human bodies—especially children, elderly individuals, and morphologies uncommon in populations. Separate models have been proposed for specific demographics, such as SMIL [hesse2018smil] for infants or distinct male/female SMPL variants, and more recent efforts like SMPL-A [agora] interpolate between children and adults. Yet, obtaining representative data for all body types remains impractical and privacy-sensitive.
@@
Figure 2: Example of local morphological changes covered by the model.
Because it is open and semantically defined, Anny can be easily inspected, extended, and integrated into existing pipelines. This design enables a wide range of applications — from millimeter-accurate scan fitting to controllable synthetic data generation and HMR — within a single unified framework. To demonstrate its potential, we create Anny-One, a large-scale synthetic dataset of 800k photorealistic humans generated with Anny, featuring expressive full-body poses, hands, and faces across diverse environments. HMR models trained with Anny-One achieve competitive accuracy on standard benchmarks and outperform existing approaches when body-shape diversity is high — using a single parametric model.
Our main contributions are: (A) Anny, a differentiable, scan-free human body model representing continuous and interpretable shape variations across genders, ages, and body types. (B) Anny-One, a large-scale synthetic dataset of 800k photorealistic humans with diverse 3D poses and shapes; (C) Empirical validation showing that Anny enables accurate scan fitting and competitive HMR, offering a unified and complete approach to human modeling. This work advocates for moving beyond scan-driven latent spaces toward open, semantically meaningful, and globally representative human models. Code will be released under the Apache 2.0 license to encourage accessible foundation tool for human-centric 3D modeling.
@@ 2 Related work
have been used to model the appearance of objects since the early days of computer graphics. Skeleton-based deformations and shape blending have been a standard way to model variability in pose and shape for decades [lewis2000posespace], and are particularly suited for human characters. Important research efforts have been made to build accurate human models representative of body surface deformations across different poses and across different individuals characteristics, through data-driven approaches based on 3D human scans. The SMPL [loper2015smpl] body model family is arguably the most widely used in the computer vision community. It relies on linear blend shapes to model shape-dependent vertex displacements in rest pose, as well as pose-corrective blend shapes to further refine the output mesh depending on the pose. The original model was extended to SMPL-X [pavlakos2019smplx] to capture full-body meshes, i.e., including facial expressions and hand poses. Other refinements were also considered, in particular STAR [osman2020star] proposed a more compact formulation compared to SMPL, and SUPR [osman2022supr] split the model into individual body parts with a sparse factorization of pose-corrective blend shapes. To model infants, a separate model SMIL [hesse2018smil] was also proposed, sometimes combined with SMPL through interpolation, referred-to as SMIL-X [agora]. Other body models with similar properties have been proposed, such as GHUM and GHUML [xu2020ghum] that use the latent space of an auto-encoder as parametrization space. More recently, ATLAS [park2025atlas] introduced a high-fidelity, scan-based model that decouples surface and skeletal representations for finer control over body proportions.
@@
A major challenge for parametric body models is to capture the diversity of body shapes observed around the world. This is hard to achieve because collecting data such as 3D body scans is costly and time consuming. Yet the above approaches, namely the SMPL and GHUM families, tackle this by learning shape representations from datasets of 3D scans. In particular, the CAESAR [robinette2002caesar] dataset used to train SMPL comprises people from the USA, Netherlands and Italy aged between 18 and 65 years in early 2000s. It contains fewer than $5000$ individuals, and already constitutes a significant data collection effort. This introduces limits on body models fitted from data. We instead rely on expert artistic manual definitions of the human body. Accurately modeling the distribution of human shape has been the focus of much efforts from computer graphics designers and artists; we leverage these efforts instead of collecting real-world data.
@@
as introduced in the eponymous work HMR [kanazawa2018end], aims to estimate a human mesh given an image. Here, we focus on parametric HMR, which refers to models that output the parameters of a parametric human body model such as SMPL [loper2015smpl]. Existing methods can be broadly separated in two categories: single-person and multi-person methods. Single-person methods assume that detections or bounding boxes for humans are provided, either from ground truth data or by an off-the-shelf detection model. In that case the HMR model directly regresses body model parameters [kanazawa2018end]. In that setting, progress was made at the level of architectures and backbones [goel2023humans, armando2024cross], or camera models [wang2023zolly, patel2025camerahmr, wang2024blade], and extensions to expressive body models [multi-hmr2024, cai2023smpler, lin2023one, sun2024aios, feng2021collaborative]. In multi-person HMR methods [multi-hmr2024, sun2021monocular, sun2024aios, sun2022putting], all humans in the input have to be detected and placed in the 3D scene by the model, in addition to regressing each human mesh. In that case, the input domain of the model typically shifts from center crops of people to images containing multiple people at arbitrary locations. Our proposed Anny-One dataset is suitable for both settings, and we empirically evaluate it with both single-person and multi-person state-of-the-art methods.
@@
Another avenue of research to improve the performance of HMR models is to focus on obtaining better training data. One of the main limiting factors in training HMR models is the lack of images labeled with 3D ground-truth humans. While earlier works [pavlakos2019smplx, bogo2016keep, huang2017towards, lassner2017unite, guler2018densepose] used optimization procedures to fit parametric models to 2D observations such as 2D keypoints, most regression-based models cited above use datasets with 2D [lin2014microsoft, andriluka14cvpr, johnson2010clustered] or 3D ground-truth joints [ionescu2013human3, mono-3dhp2017], sometimes with pseudo-ground-truth meshes [li2022cliff, joo2021exemplar, kolotouros2019learning, moon2022neuralannot, lin2023one]. Acquiring images with 3D ground truth is expensive, time-consuming, and pseudo-annotations often lack precision in particular for hands and faces. A popular solution is to use synthetic data; it presents 3 major advantages: 1) scaling is efficient 2) annotations are free of noise 3) the diversity of the data, in terms of many variables such as body shapes, skin tones, ages, motions, environments, or cameras is easy to control. Existing works have shown that synthetic data can complement real data [varol2017learning, agora], and more recently that synthetic data alone leads to state-of-the-art (SOTA) HMR models [bedlam, yin2024whac, multi-hmr2024], in particular when expressive poses (including hands and faces) are involved [multi-hmr2024] or accurate camera estimation in human-centric scenes is required [wang2024blade, patel2025camerahmr].
In this work, we leverage all advantages of synthetic data. Our large-scale Anny-One dataset contains diverse humans in terms of body shape and appearance, with expressive poses including faces and hands, with exact ground truth. We train expressive single-person and multi-person HMR models, and obtain SOTA performance without using real images for training.
@@ 3 Modeling Anny Body
We propose Anny, a differentiable parametric mesh model aiming at modeling a large diversity of human morphologies. Anny is built upon assets from MakeHuman [makehuman, mpfb2], a community-driven framework available under an open-source license that enables artists to model a wide variety of human-like 3D characters. We release a PyTorch implementation of Anny under a permissive Apache 2.0 open-source license to foster the development of human-centric research and applications.
@@
The base model of Anny is composed of 13,380 vertices and 13,378 quadrilateral faces (excluding tongue and eyes), assigned with weights to a default rig of 163 bones, illustrated Fig. 3. Other rigs are additionally supported, such as the one typically used in Mixamo [mixamo] character animations, which allows for simple animation retargeting by reproducing bone orientations.
Figure 3: Skeleton of Anny. Left: default skeleton shown for different morphologies, excluding facial bones for visualization. Right: use of the same skeleton as a Mixamo [mixamo] character animation for retargeting purposes.
@@
Following MakeHuman terminology, we capture the diversity of human morphology through various parameters, which we refer to as phenotypes. These phenotypes aim to encode high-level characteristics, such as age (Fig. 1), gender, weight, muscle amount, etc., as well as more local changes, such as the amount of head fat or changes in belly morphology during pregnancy (Fig. 2). Phenotypes are grounded into a set of prototypical blendshapes, each modeling a 3D human mesh with particular characteristics, e.g. a female baby having small muscle, average weight, and large height.
@@
Phenotypes are based on preconceptions of artists regarding particular human traits. As a result, they encode by design stereotypes of MakeHuman artists, and one should not expect phenotype parameters to faithfully encode any identity-related characteristics, such as gender, age or ethnicity. We nonetheless found these parameters useful to model the diversity of human morphologies, and we chose to keep MakeHuman terminology to make explicit the intent of the artists. Existing data-driven models [loper2015smpl, pavlakos2019smplx, osman2020star, xu2020ghum, osman2022supr, park2025atlas] avoid addressing this delicate question by using abstract shape spaces with no explicit semantics–except for gender-specific models. They nonetheless also convey biases related to their topology (all mesh models, including Anny, assume individuals with four limbs) and to their training distribution (most models are only suited for adults, for example).
Figure 4: Model calibration. Calibration of Anny morphological shape distribution with WHO Child Growth standards for boys [who2006growth].
@@
Anny takes shape and pose parameters as input and outputs a 3D posed human mesh representation in a backward-differentiable manner. The shape of the mesh is controlled by scalar coefficients targeting various phenotypes regarding age, weight, gender, etc.. These coefficients define weights of a piecewise-multilinear interpolation between prototypical blendshapes, that are used to adjust the shape of a base model. For instance, blendshapes corresponding to the phenotypes child (age=1/3age{=}1/3) and young (age=2/3age{=}2/3) contribute equally to the mesh deformation given a parameter value age=0.5age{=}0.5, as illustrated in Fig. 1. This use of interpolation constraints the shape space structure and helps producing topologically consistent shape deformations. The shape-adjusted model features a skeleton, composed of a set of bones connected along a kinematic tree. One bone is defined as root bone of the tree. To encode a pose for the model, the pose of the root bone is given, together with 3D rotations at each joint between connected bones, relative to their rest configuration. Given pose parameters, we apply forward kinematics to retrieve bone poses. We then deform the mesh using blend skinning, producing the final human mesh, shaped and posed accurately according to the input parameters. These steps are implemented using PyTorch [paszke2017automatic] and NVIDIA Warp [warp2022] to benefit from the automatic gradient backpropagation features of these libraries.
@@ 4 Shape statistical modeling
Having a model of the distribution of human morphologies is useful for applications such as data synthesis, or as an optimization prior. Existing body models [loper2015smpl, park2025atlas] typically feature a latent space that encodes the shape distribution from their training set, e.g. using PCA for SMPL. These training sets consist of proprietary collections of 3D scans which required significant acquisition effort, yet they remain insufficient to represent the global population. Datasets such as CAESAR [robinette2002caesar] and SizeUSA [sizeusa] consist of scans of adults from western countries, and as such are not representative of morphological variations among all ages and all global population. In contrast, Anny can model a large diversity of morphologies thanks to its various phenotype parameters. This diversity covers common body shapes but also uncommon ones, such as 2.4 meters high individuals. We therefore calibrate the distribution of Anny parameters to obtain a shape distribution more resembling of the global population. We empirically define a bijective mapping between the age parameter of Anny and some morphological age in years. We then model distributions for major phenotype parameters as Beta distributions, conditioned on age and gender. These distributions are jointly calibrated to match the mean and standard deviation of reference growth standards and body mass index from the World Health Organization (WHO) [who2006growth], as shown in Figure 4.

### v2

@@ Abstract
Parametric body models provide the structural basis for many human-centric tasks, yet existing models often rely on costly 3D scans and learned shape spaces that are proprietary and demographically narrow. We introduce Anny, a simple, fully differentiable, and scan-free human body model grounded in anthropometric knowledge from the MakeHuman community. Anny defines a continuous, interpretable shape space, where phenotype parameters (e.g. gender, age, height, weight) control blendshapes spanning a wide range of human forms–across ages (from infants to elders), body types, and proportions. Calibrated using WHO population statistics, it provides realistic and demographically grounded human shape variation within a single unified model. Thanks to its openness and semantic control, Anny serves as a versatile foundation for 3D human modeling–supporting millimeter-accurate scan fitting, controlled synthetic data generation, and Human Mesh Recovery (HMR). We further introduce Anny-One, a collection of 800k photorealistic images generated with Anny, showing that despite its simplicity, HMR models trained with Anny can match the performance of those trained with scan-based body models. The Anny body model and its code are released under the Apache 2.0 license, making Anny an accessible foundation for human-centric 3D modeling.
@@ 1 Introduction
Understanding human shape and motion in 3D is fundamental to computer vision, graphics, and robotics. Parametric body models provide a compact and differentiable representation of the human body, enabling a wide range of applications–from scan fitting and animation to Human Mesh Recovery (HMR) from images and videos [goel2023humans, yin2025smplest, zhang2021pymaf, shin2024wham, kocabas2020vibe]. These models have been central to progress in human-centric perception, powering tasks such as motion analysis, behavior understanding, and human–robot interaction.
Among them, the SMPL family [loper2015smpl, pavlakos2019smplx, osman2020star, osman2022supr] has played a transformative role. SMPL-like models achieve high accuracy on standard benchmarks by learning low-dimensional shape spaces from 3D body scans [robinette2002caesar, Bogo:CVPR:2014, Bulat2017FAW]. However, this data-driven design also imposes limitations: the available scan datasets are expensive, demographically narrow, and privacy-sensitive. As a result, current models struggle to represent the full diversity of human bodies—especially children, elderly individuals, and morphologies uncommon in populations. Separate models have been proposed for specific demographics, such as SMIL [hesse2018smil] for infants or distinct male/female SMPL variants, and more recent efforts like SMPL+A [agora] interpolate between children and adults. Yet, obtaining representative data for all body types remains impractical and privacy-sensitive.
@@
Because it is open and semantically defined, Anny can be easily integrated into existing pipelines. This design enables a wide range of applications — from millimeter-accurate scan fitting to controllable synthetic data generation and HMR — within a single unified framework. To demonstrate its potential, we create Anny-One, a large-scale synthetic dataset of 800k photorealistic humans generated with Anny, featuring expressive full-body poses, hands, and faces across diverse environments. HMR models trained with Anny-One achieve competitive accuracy on standard benchmarks and outperform existing approaches when body-shape diversity is high — using a single parametric model.
Our main contributions are: (A) Anny, a differentiable, scan-free human body model representing continuous and interpretable shape variations across genders, ages, and body types. (B) Anny-One, a large-scale synthetic dataset of 800k photorealistic humans with diverse 3D poses and shapes; (C) Empirical validation showing that Anny enables accurate scan fitting and competitive HMR, offering a unified and complete approach to human modeling. This work advocates for moving beyond scan-driven latent spaces toward open, semantically meaningful, and globally representative human models. Code will be released under the Apache 2.0 license to serve as an accessible foundation tool for human-centric 3D modeling.
@@
Figure 2: Example of local morphological variations covered by the model.
@@ 2 Related work
have been used to model the appearance of objects since the early days of computer graphics. Skeleton-based deformations and shape blending have been a standard way to model variability in pose and shape for decades [lewis2000posespace], and are particularly suited for human characters. Important research efforts have been made to build accurate human models representative of body surface deformations across different poses and across different individuals characteristics, through data-driven approaches based on 3D human scans. The SMPL [loper2015smpl] body model family is arguably the most widely used in the computer vision community. It relies on linear blend shapes to model shape-dependent vertex displacements in rest pose, as well as pose-corrective blend shapes to further refine the output mesh depending on the pose. The original model was extended to SMPL-X [pavlakos2019smplx] to capture full-body meshes, i.e., including facial expressions and hand poses. Other refinements were also considered, in particular STAR [osman2020star] proposed a more compact formulation compared to SMPL, and SUPR [osman2022supr] split the model into individual body parts with a sparse factorization of pose-corrective blend shapes. To model infants, a separate model SMIL [hesse2018smil] was also proposed, sometimes combined with SMPL-X through interpolation [agora], referred-to as SMPL-X+A. Other body models with similar properties have been proposed, such as GHUM and GHUML [xu2020ghum] that use the latent space of an auto-encoder as parametrization space. More recently, ATLAS [park2025atlas] introduced a high-fidelity, scan-based model that decouples surface and skeletal representations for finer control over body proportions.
@@
A major challenge for parametric body models is to capture the diversity of body shapes observed around the world. This is hard to achieve because collecting data such as 3D body scans is costly and time consuming. Yet the above approaches, namely the SMPL and GHUM families, tackle this by learning shape representations from datasets of 3D scans. In particular, the CAESAR [robinette2002caesar] dataset used to train SMPL comprises people from the USA, Netherlands and Italy aged between 18 and 65 years in early 2000s. It contains fewer than $5000$ individuals, and already constitutes a significant data collection effort. This introduces limits on body models fitted from scan data. We instead rely on expert artistic manual definitions of the human body. Accurately modeling the variability of human shape has been the focus of much efforts from computer graphics designers and artists; we leverage these efforts instead of collecting real-world data.
@@
as introduced in the eponymous work HMR [kanazawa2018end], aims to estimate a human mesh given an image. Here, we focus on parametric HMR, which refers to models that output the parameters of a parametric human body model such as SMPL [loper2015smpl]. Existing methods can be broadly separated in two categories: single-person and multi-person methods. Single-person methods assume that human detections or bounding boxes are provided, obtained either from ground truth data or from an off-the-shelf detection model. In that setting, the HMR model directly regresses body model parameters [kanazawa2018end]. Progress was made at the level of architectures and backbones [goel2023humans, armando2024cross], camera models [wang2023zolly, patel2025camerahmr, wang2024blade], and extensions to expressive body models [multi-hmr2024, cai2023smpler, lin2023one, sun2024aios, feng2021collaborative]. In multi-person HMR [multi-hmr2024, sun2021monocular, sun2024aios, sun2022putting], all visible humans in the input have to be detected and localized in 3D by the model, in addition to regressing each human mesh. In that case, the input domain of the model typically shifts from center crops of people to images containing multiple people at arbitrary locations. Our proposed Anny-One dataset is suitable for both settings, and we empirically evaluate it with both single-person and multi-person state-of-the-art methods.
@@
Another avenue of research to improve the performance of HMR models is to focus on obtaining better training data. One of the main limiting factors in training HMR models is the lack of images labeled with 3D ground-truth humans. While earlier works [pavlakos2019smplx, bogo2016keep, huang2017towards, lassner2017unite, guler2018densepose] used optimization procedures to fit parametric models to 2D observations such as 2D keypoints, most regression-based models cited above use datasets with 2D [lin2014microsoft, andriluka14cvpr, johnson2010clustered] or 3D ground-truth joints [ionescu2013human3, mono-3dhp2017], sometimes with pseudo-ground-truth meshes [li2022cliff, joo2021exemplar, kolotouros2019learning, moon2022neuralannot, lin2023one]. Acquiring images with 3D ground truth is expensive, time-consuming, and pseudo-annotations often lack precision in particular for hands and faces. A popular solution is to use synthetic data; it presents 3 major advantages: 1) scaling is efficient 2) annotations are free of noise 3) the diversity of data, in terms of many variables such as body shapes, skin tones, ages, motions, environments, or cameras is easy to control. Existing works have shown that synthetic data can complement real data [varol2017learning, agora], and more recently that synthetic data alone leads to state-of-the-art (SOTA) HMR models [bedlam, yin2024whac, multi-hmr2024], in particular when expressive poses (including hands and faces) are involved [multi-hmr2024] or accurate camera estimation in human-centric scenes is required [wang2024blade, patel2025camerahmr]. In this work, we leverage all advantages of synthetic data. Our large-scale Anny-One dataset contains diverse humans in terms of body shape and appearance, with expressive poses including faces and hands, with exact ground truth. We train expressive single-person and multi-person HMR models, and obtain SOTA performance without using real images for training.
@@ 3 Modeling Anny Body
We propose Anny, a differentiable parametric mesh model aiming at modeling a large diversity of human morphologies. Anny is built on assets from MakeHuman [makehuman, mpfb2], a free and community-driven framework that enables artists to model a wide variety of human-like 3D characters. We release a PyTorch implementation of Anny under a permissive Apache 2.0 open-source license to foster the development of human-centric research and applications.
@@
The base mesh of Anny is composed of 13,718 vertices and 13,710 quadrilateral faces (including tongue and eyes), assigned with weights to a default rig of 163 bones, illustrated Fig. 3. Anny also supports coarser mesh resolutions and additional rigs such as the one used for Mixamo characters [mixamo], which allows simple animation re-targeting by reproducing bone orientations.
Figure 3: Mesh and skeleton of Anny. Left: default mesh and skeleton, modeling two different morphologies. Middle: coarser mesh (1,229 vertices) using only a skeleton subset. Right: use of the same skeleton as a Mixamo [mixamo] character for pose re-targeting.
@@
We capture the diversity of human morphology through various parameters which we refer to as phenotypes, following MakeHuman terminology. These phenotypes aim to encode high-level characteristics, such as age (Fig. 1), gender, weight, muscle amount, etc., as well as more local changes, such as the amount of head fat or changes in belly morphology during pregnancy (Fig. 2). Phenotypes are grounded into a set of prototypical blendshapes, each modeling a 3D human mesh with particular characteristics, e.g. a female baby having small muscle, average weight, and large height.
@@
Phenotypes are based on preconceptions of artists regarding particular human traits. As a result, they encode by design stereotypes of MakeHuman artists, and one should not expect phenotype parameters to faithfully encode any identity-related characteristics, such as gender, age or ethnicity. We nonetheless found these parameters useful to model the diversity of human morphologies, and we chose to keep MakeHuman terminology to make explicit the artists’ intent. Existing scan-based models [loper2015smpl, pavlakos2019smplx, osman2020star, xu2020ghum, osman2022supr, park2025atlas] avoid addressing this delicate question by relying on abstract shape spaces with no explicit semantics – except for gender-specific models. It is worth mentioning that these models also convey biases, related to their topology (all mesh models, including Anny, assume individuals with four limbs) and to their training distribution (most models are only suited for adults, for example).
Figure 4: Model calibration. Calibration of Anny morphological shape distribution with WHO Child Growth standards [who2006growth] (boys).
@@
Anny takes shape and pose parameters as input and outputs a 3D posed human mesh representation in a backward-differentiable manner. The shape of the mesh is controlled by scalar coefficients targeting various phenotypes regarding age, weight, gender, etc.. These coefficients define weights of a piecewise-multilinear interpolation between prototypical blendshapes, that are used to adjust the shape of a base model. For instance, blendshapes corresponding to the phenotypes child (age=1/3age{=}1/3) and young (age=2/3age{=}2/3) contribute equally to the mesh deformation given a parameter value age=0.5age{=}0.5, as illustrated in Fig. 1. This use of interpolation constraints the structure of the shape space and helps producing topologically consistent meshes. The shape-adjusted model features a skeleton, composed of a set of bones connected along a kinematic tree. One bone is defined as root bone of the tree. To encode a pose for the model, the pose of the root bone is given, together with 3D rotations at each joint between connected bones, relative to their rest configuration. Given pose parameters, we apply forward kinematics to retrieve bone poses. We then deform the mesh using blend skinning, producing the final human mesh, shaped and posed accurately according to the input parameters. These steps are implemented using PyTorch [paszke2017automatic] and NVIDIA Warp [warp2022] to benefit from the automatic gradient backpropagation features of these libraries.
@@ 4 Shape statistical modeling
Having a model of the distribution of human morphologies is useful for applications such as data synthesis, or as an optimization prior. Existing body models [loper2015smpl, park2025atlas] typically feature a latent space that encodes a shape distribution derived from their training sets, e.g. using PCA for SMPL. These training sets consist of proprietary collections of 3D scans which required significant acquisition effort, yet they remain insufficient to represent the global population. Datasets such as CAESAR [robinette2002caesar] and SizeUSA [sizeusa] consist of scans of adults from western countries, and as such are not representative of morphological variations among the global population and all ages. In contrast, Anny can model a large diversity of morphologies thanks to its various phenotype parameters. This diversity covers common body shapes but also uncommon ones, such as 2.4 meters high individuals. We therefore calibrate the distribution of Anny parameters to obtain a shape distribution more resembling of the global population. We empirically define a bijective mapping between the age parameter of Anny and some morphological age in years. We then model distributions for major phenotype parameters as Beta distributions, conditioned on age and gender. These distributions are jointly calibrated to match the mean and standard deviation of reference growth standards and body mass index from the World Health Organization (WHO) [who2006growth], see Figure 4.